OpenStudy (anonymous):

In a recent school election 980 votes were cast. The winner recieved 372 votes more than the loser. How many votes did each of the two candidates recieve.

OpenStudy (anonymous):

Just set the total equal to the number of loser votes plus the number of loser votes + 372, then solve for x 980 = x + (x+372) 980 = 2x + 372

OpenStudy (anonymous):

x + y = 980 x = y +372 980 = 2y +372 y = 304 ; x = 676